OpenStudy (anonymous):

A geostationary satellite is positioned 35,800 km above Earth’s surface. It takes 24 hours to complete one orbit. The radius of Earth is about 6400 km. a) What distance does the satellite travel in 1 h? 3 h? 2.5 h? 25 h? b) After how many hours has the satellite traveled 200,00 km?

OpenStudy (jack1):



OpenStudy (anonymous):

That is literally the only part I know how to figure out. :P

OpenStudy (anonymous):

Would I just find the circumference of the orbit and then divide into chunks based on the time intervals??

OpenStudy (jack1):

perfect, that solution sounds really good

OpenStudy (jack1):

compare answers here?

OpenStudy (anonymous):

Circumference = approx. 265,016 so, 1 hr = 11,042.34 km? 3 hr = 33,127? 2.5 hr = 27,605.84? 25 hr = 276,058.34? I think this is correct, but how would I calcualate how long it took to travel 200,000 km?

OpenStudy (anonymous):

*calculate

OpenStudy (anonymous):

wait... I would just divide the circumference by 200,000 to find that... huh?

OpenStudy (jack1):

yep, what percentage of the circumference is 200000?

OpenStudy (anonymous):

75% so, 18 hrs?

OpenStudy (jack1):

so flip ur idea and reverse it divide (200000/ circumference) *100 = % of circumference

OpenStudy (jack1):

yep perfect dude

OpenStudy (anonymous):

well, like 18.1 because the percent is not perfect. :P

OpenStudy (jack1):

I got: circumference = 265150 km so it travels 265150km every 24 hours therefore 11048 km per hour so 3 hours = 3*11048 = 33134 km 2.5 hrs = 27619.8 km 25 hrs = 276198 and 75.43 % so 18 hrs and 6 minutes

OpenStudy (jack1):

but i recon they'll let the rounding slide in this case, sig figs an all that

OpenStudy (anonymous):

Sweet, I feel a little less stupid. Thanks for working with me on this! :)

OpenStudy (jack1):

always welcome dude

OpenStudy (anonymous):

Also, do you know how to use cosine functions and such? If you have a minute, I am going to post another problem. I really don't understand them. Could you help me?

OpenStudy (jack1):

dude, im sorry, is 3am here and i'm beat, so i'm gonna bail on you, sorry again, need zzzz's tho hit up @ganeshie8 , he's my goto mathlete

OpenStudy (anonymous):

haha, okay. Well thank you anyways! Have a good night!! :)

ganeshie8 (ganeshie8):

sorry WTFork im going out for a while... cya

OpenStudy (anonymous):

No worries! Thanks anyways! :)